OpenStudy (anonymous):

Christina Hercher borrowed $50,000 on a 90 day, eight percent note. Christina paid $3,000 toward the note on day 40. On day 60 she paid an additional $4,000. Using the U.S. Rule, Christina's adjusted balance after the first payment is:

OpenStudy (anonymous):

The US rule allows the borrower to receive proper interest credits. This rule states that any partial loan payment first covers any interest that has built up. The remainder of the partial payment reduces the loan principal.

OpenStudy (anonymous):

Is this answer correct? 50000x.08x90/360= 100000

OpenStudy (anonymous):

100000-3000=70000

OpenStudy (anonymous):

they want to know what the adjusted balance is of the FIRST payment. so its: 50,000 x .08 x 40/360 = 444.44